OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + x2 – 6x

jhonyy9 (jhonyy9):

one solution

jhonyy9 (jhonyy9):

0,

jhonyy9 (jhonyy9):

2

OpenStudy (anonymous):

x(x^2+x-6) 1st x=0, obviously. x^2+x-6=0 x=-0,5+-sqrt(0,5^2+6) x=-0,5+-2,5 2nd x=-0,5+2,5=2 3rd x=-0,5-2,5=-3

OpenStudy (anonymous):

that answer is not one of the multiple choice answers that they give

OpenStudy (anonymous):

0, –2, –3 0, –2, 3 0, 2, –3 0, 2, 3

OpenStudy (anonymous):

these are the answers they give to choose from

OpenStudy (anonymous):

0, 2, -3 is what I said?

OpenStudy (anonymous):

Well, meant atleast...

OpenStudy (anonymous):

awesome thankx